Find the probability of throwing 11 each time in 3 tosses of 2 dice.
1
Expert's answer
2020-03-05T09:38:36-0500

2 dice can throw 11 in two cases: 5 : 6 or 6 : 5.

The probability to give any specific number for 1 dice is 1/6. For 2 dice -- 1/6 * 1/6 = 1/36. But as there are two ways to throw 11, then the probability to throw 11 in one toss is 2*1/36 = 1/18

As we have 3 tosses than the probability is


"(1\/18)^3 = 1\/5832 = 0.00017"

Answer: the probability of throwing 11 each time in 3 tosses of 2 dice is 0.00017